a bullet moving with a velocity of 200 cm per second penetrates wooden block and comes to rest after transferring 4 cm inside 8 what velocity is needed for travelling distance of 9 CM in same block

Grade:11

2 Answers

Vikas TU
6 years ago
Dear Student,
Given Velocity = 200 cm/sec
Initial velocity = 0
From the Netwon's third Law of equation
V^2- U^2 = 2as
40000 = 8 a
a = 5000 cm/sec^2
v^2 = 40000+ 90000
             V^2 = 130000
           V = 100√13 cm/sec
This would be the velocity.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Bhumika
5 years ago
V=0 ; u=200 cm/sec ; distance =s=4 cm
From the third equation of motion:
V^2-u^2= 2as
0-(200)^2=2a.4
40000=8a
a=5000cm/sec^2
To travel 9 cm in a wooden block:
U=0;distance =9 cm
From the third equation of motion
V^2-u^2=2as
V^2-0=2×5000×9
V=300cm/sec